# Newtonian potential from scattering amplitudes in super-renormalizable gravity

### v1

@@ Abstract
In calculations, it is convenience to separate tensors into parallel and orthogonal components relative to the momentum $k_{\mu}$ of the scalar particle by using the projection operators: $\eta^{||}_{\mu\nu}=\frac{k_{\mu}k_{\nu}}{m^{2}}$, and $\eta^{\bot}_{\mu\nu}=n_{\mu\nu}-\frac{k_{\mu}k_{\nu}}{m^{2}}$, which are diagonal in the inertial frame of $k_{\mu}$ and represent the time and spatial components of $\eta_{m\nu}$, respectively.
In the classical limit, contributions from Feynman diagrams persist to all loop orders, with each additional loop introducing higher powers of the gravitational coupling $G_{N}$. Physically, the long-range interaction is dominated by the region where graviton momentum $q\to 0$, corresponding to a heavy scalar source of mass $m$ at rest. As depicted in Fig. 1, the amplitude in this regime encodes the metric perturbation $h_{\mu\nu}$ generated by the point-like source and its gravitational field. This amplitude, denoted $\mathcal{M}^{\mu\nu}_{\text{vertex}}$, incorporates both the energy-momentum pseudotensor and gauge-fixing contributions. A relation derived in Jakobsen:2020ksu links the gauge-invariant part of the scattering amplitude to the physical pseudotensor as
@@
where $\tilde{\tau}^{\mu\nu}$ represents the combined contribution from the matter ($T_{\mu\nu}$) and the nonlinear gravitational field. To zeroth order, it corresponds to the energy-momentum tensor of point particle in special relativity, while loop corrections account for the energy-momentum contribution from the surrounding, self-interacting gravitational field. The term $\tilde{H}^{\mu\nu}_{\text{non-linear}}$ arises from the nonlinear part of the gauge-fixing term is chosen as: $S_{\text{gf}}=\int\mathrm{d}^{D}x\,\eta^{\mu\nu}\partial_{\alpha}h^{\alpha}_{\mu}\,\omega(\Box)\,\partial_{\beta}h^{\beta}_{\nu}/(\kappa^{2}\xi)$. Due to the conservation of momentum in both the initial and final states, the constraint $\delta(kq)=\delta(mq_{||})$ must be imposed to properly define the integral in classical limit. To obtain the correction to the metric, we perform a Fourier transform of the above expression and combine it with the equations of motion, resulting in
@@
$g_{\mu\nu}=\eta_{\mu\nu}-\frac{\kappa}{2}\int\frac{\,\mathrm{d}^{D}q\delta(kq)e^{-iqx}}{(2\pi)^{D-1}}O^{-1}_{\alpha\beta\mu\nu}\mathcal{M}_{\mathrm{vertex}}^{\alpha\beta},$ (5)
where $O^{-1}_{\alpha\beta\mu\nu}$ represents the propagator in gravitational theory. For GR, it takes the form $O^{-1}_{\alpha\beta\mu\nu}=\frac{I_{\alpha\beta\mu\nu}-\frac{1}{D-2}\eta_{\mu\nu}\eta_{\alpha\beta}}{q^{2}+i\epsilon}$, corresponding to the Schwarzschild-Tangherlini solution Jakobsen:2020ksu . For the sake of simplifying calculations, we examine a particular class of form factors: $\omega(\Box)=e^{-\Box/M^{2}_{\ast}}$, and $H_{0}(\Box)=H_{2}(\Box)=-\Box/M^{2}_{\ast}$. Therefore $F_{1}(\Box)=-F_{2}(\Box)/2=\left(e^{-\Box/M^{2}_{\ast}}-1\right)/\Box$. It is straightforward to show that these form factors simplify the propagator as
$\begin{split}O^{-1}_{\mu\nu\alpha\beta}&=\frac{e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}I_{\mu\nu\alpha\beta}-\frac{4e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{(D-2)q^{2}}T_{\mu\nu\alpha\beta}+\frac{2e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{(D-2)q^{2}}C_{\mu\nu\alpha\beta}\\ &+\frac{(2\xi-2)e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}J_{\mu\nu\alpha\beta}+\frac{e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}\left(\frac{D-2}{D-1}-\frac{1}{(D-2)(D-1)}-\frac{3\xi}{2}\right)K_{\mu\nu\alpha\beta},\end{split}$ (6)
@@
where the five independent operators are $I_{\alpha\beta}^{\mu\nu}=\frac{1}{2}\left(\delta_{\alpha}^{\mu}\delta_{\beta}^{\nu}+\delta_{\beta}^{\mu}\delta_{\alpha}^{\nu}\right)$, $T_{\alpha\beta}^{\mu\nu}\equiv\frac{1}{4}\eta^{\mu\nu}\eta_{\alpha\beta}$, $C_{\alpha\beta}^{\mu\nu}\equiv\frac{1}{2}\left(\eta^{\mu\nu}\frac{q_{\alpha}q_{\beta}}{q^{2}}+\frac{q^{\mu}q^{\nu}}{q^{2}}\eta_{\alpha\beta}\right)$, $J_{\alpha\beta}^{\mu\nu}\equiv I_{\rho\kappa}^{\mu\nu}\frac{q_{\sigma}q^{\rho}}{q^{2}}I_{\alpha\beta}^{\sigma\kappa}$, and $K_{\alpha\beta}^{\mu\nu}\equiv\frac{q^{\mu}q^{\nu}}{q^{2}}\frac{q_{\alpha}q_{\beta}}{q^{2}}$. With the Feynman rules, the tree-level scattering amplitude is given by $\mathcal{M}^{\mu\nu}_{\text{tree}}=-\kappa m^{2}\eta^{||}_{\mu\nu}$, reflecting the leading coupling of massive scalars to gravity. Based on the coupling between the propagator and the scattering amplitudes (5), we calculate the first-order correction to the metric tensor as
@@
In particular, for $D=4$, the Newtonian potential reduces to $\Phi^{(G_{N})}(r)=-\frac{G_{N}m}{r}\text{erf}\left(\frac{M_{\ast}r}{2}\right)$, where the error function is defined as $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}dt$. This result aligns with earlier work in Biswas:2011ar obtained with linear perturbation. The error function encodes the nonlocal smearing of the gravitational interaction: at large distances $\lim_{x\to\infty}\text{erf}(x)=1$, the potential reduces to the standard Newtonian form $\frac{G_{N}m}{r}$, recovering GR in the infrared. In the short-range regime, however, the potential remains perfectly regular: $\Phi^{(G_{N})}(r\to 0)=-\frac{G_{N}mM_{*}}{\sqrt{\pi}}$. The weak nonlocality regulates the ultraviolet behavior of gravity, defusing the classical singularity at tree level.
Newtonian potential at one-loop level. At order $(G_{N})^{2}$, the classical metric correction originates from the triangle one-loop diagram (Fig. 2), which uniquely provides the non-analytic terms responsible for long-range interactions Bjerrum-Bohr:2018xdl . Other one-loop topologies, devoid of such non-analyticity, do not contribute to the classical potential, thereby establishing the triangle diagram as the sole contributor to the $\mathcal{O}(G_{N}^{2})$ singularity resolution.
@@
where $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}$ and $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}$ denote the three-graviton vertex interactions originating from the Einstein–Hilbert term and the nonlocal sector of the action, respectively. A key element in the amplitude is the tensor
$f_{\alpha\beta}\equiv\frac{D-3}{D-2}\eta^{\parallel}_{\alpha\beta}-\frac{1}{D-2}\eta^{\perp}_{\alpha\beta},$ (10)
which effectively encodes the scalar–graviton vertices contracted with the graviton propagator. This tensor captures the physical structure inherited from the $\phi^{2}h$ coupling after integrating out intermediate graviton modes.
In the classical limit, where the graviton momenta are soft compared to the momentum of the massive scalar particle, the scalar propagator admits a significant simplification. Specifically, terms involving graviton momenta $l$ and $q$ in the scalar–graviton vertex, as well as the momentum-dependent tensor structures in the propagator numerator, become subdominant and can be systematically neglected. This soft-graviton approximation isolates the leading long-range interaction and is essential for extracting the classical dynamics from the amplitude.
Considering on the tensor part of the integral involving the three-graviton vertex from the GR contribution $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}\right)$ and nonlocal part $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}\right)$, yields
@@
where $U$-tensor is defined in Appendix A, and the quantities $I^{\text{GR}}_{a_{1}...a_{n}}$, and $I^{(i)}_{a_{1}...a_{n}}$ ($i=1,2,3$) are given by
$\begin{split}&I^{\text{GR}}_{a_{1}...a_{n}}\equiv\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}l_{a_{1}}...l_{a_{n}}}{l^{2}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (12)
$\begin{split}&I^{(1)}_{a_{1}...a_{n}}\equiv-\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}-1\right)l_{a_{1}}...l_{a_{n}}}{l^{2}(l+q_{\bot})^{4}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (13)
$\begin{split}&I^{(2)}_{a_{1}...a_{n}}\equiv-\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}-1\right)l_{a_{1}}...l_{a_{n}}}{l^{2}q^{2}_{\bot}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)},\end{split}$ (14)
$\begin{split}&I^{(3)}_{a_{1}...a_{n}}\equiv-\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}\left(e^{\frac{l^{2}}{M^{2}_{\ast}}}-1\right)l_{a_{1}}...l_{a_{n}}}{l^{4}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)}.\end{split}$ (15)
@@
Ultimately, the corrections to the Newtonian potential up to second order in $(G_{N})^{2}$ within the one-loop approximation can be expressed as
@@
$\begin{split}\Phi^{(G_{N})^{2}}(r)&=-\frac{\kappa^{4}m^{2}}{4}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}e^{-iq_{\bot}x_{\bot}}e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}\left[N_{D-1}\frac{183D+5D^{3}-61D^{2}-151}{16(D-2)^{2}}\right.\\ &\left.-q^{2}_{\bot}\left(N^{(1)}_{D-1}+N^{(3)}_{D-1}\right)\left(\frac{D^{2}-5D+8}{4(D-2)^{3}}\right)-q^{2}_{\bot}N^{(2)}_{D-1}\left(\frac{76+13D+10D^{2}}{8(D-2)^{3}}\right)\right],\end{split}$ (16)
@@
where the integrals $N_{D-1},N^{(i)}_{D-1},i=1,2,3$ are given in Appendix B. The Newtonian potential at order $\mathcal{O}(G_{N}^{2})$, given in Eq. (16), is structurally governed by the spacetime dimension $D$ and four loop integrals $N_{D-1},N^{(i)}_{D-1}$ (with $i=1,2,3$), whose explicit forms are provided in Appendix B. These integrals do not admit a simple analytic expression, motivating a numerical treatment in the physically relevant case $D=4$. As displayed in Fig. 3, the computed potential is plotted against the dimensionless radial coordinate, the one-loop correction grows smoothly with distance, matching expected long-range behavior. Most notably, the potential remains completely regular as $r\to 0$, offering direct numerical evidence that the mechanism of singularity resolution, enforced by weak nonlocality, persists at the quantum level.

### v2

@@ III The Metric correction from Tree-Level Scattering Amplitudes
where the scalar fields are treated as neutral, non-rotating point particles. To simplify, we represent the matter and gravitational field using minimal coupling. Then, we proceed by expanding the action in terms of powers of the fields
$S_{\text{matter}}$ $\approx\frac{1}{2}\int d^{D}x\left(\eta^{\mu\nu}\phi_{,\mu}\phi_{,\nu}-m^{2}\phi^{2}\right)$ (14)
$-\frac{1}{2}\int d^{D}x\left(h^{\mu\nu}\phi_{,\mu}\phi_{,\nu}-\frac{1}{2}h_{\mu}^{\mu}\left(\phi^{,\nu}\phi_{,\nu}-m^{2}\phi^{2}\right)\right).$
From the $\phi^{2}h$ action we get tree-lever vertex rule proportional to $i\kappa V^{\mu\nu}_{\text{vertex}}$, where
$V_{\text{vertex}}^{\mu\nu}(k,k-q,m)\equiv I_{\alpha\beta}^{\mu\nu}k^{\alpha}(k-q)^{\beta}-\frac{k(k-q)-m^{2}}{2}\eta^{\mu\nu}$ (15)
@@
Moreover, the Lorentz invariance of the perturbative QFT permits the use of an arbitrary inertial reference frame. To streamline the analysis, we introduce a notation that separates tensors into parallel and orthogonal components relative to the momentum $k_{\mu}$ of the scalar particle. This leads to the definition of the following projection operators
$\left\{\begin{split}&\eta^{||}_{\mu\nu}=\frac{k^{\mu}k^{\nu}}{m^{2}},\\ &\eta^{\bot}_{\mu\nu}=n_{\mu\nu}-\frac{k^{\mu}k^{\nu}}{m^{2}}.\end{split}\right$. (16)
These two operators are independent, and their sum is equal to $\eta_{\mu\nu}$. In the inertial frame of $k_{\mu}$, these components simplify significantly, becoming diagonal and representing the time and spatial components, respectively. As depicted in Fig. 1, the amplitude in this regime encodes the metric perturbation $h_{\mu\nu}$ generated by the point-like source and its gravitational field. This amplitude, denoted $\mathcal{M}^{\mu\nu}_{\text{vertex}}$, incorporates both the energy-momentum pseudotensor and gauge-fixing contributions. A relation derived in Jakobsen (2020) links the gauge-invariant part of the scattering amplitude to the physical pseudotensor as
@@
where $\tilde{\tau}^{\mu\nu}$ represents the combined contribution from the matter ($T_{\mu\nu}$) and the nonlinear gravitational field. To zeroth order, it corresponds to the energy-momentum tensor of point particle in special relativity, while loop corrections account for the energy-momentum contribution from the surrounding, self-interacting gravitational field. $\tilde{H}^{\mu\nu}_{\text{non-linear}}$ denotes the total contribution from the nonlinear gauge-fixing term. However, we adopt Eq.(6) as the gauge-fixing term, which does not contribute any nonlinear terms. Therefore, this term can be safely neglected in our calculations. The linear term appears as a dynamical propagator, which contributes to Eq.(8). Furthermore, due to the conservation of momentum in both the initial and final states, the constraint $\delta(kq)=\delta(mq_{||})$ must be imposed to properly define the integral in classical limit. To obtain the correction to the metric, we perform a Fourier transform of the above expression and combine it with the equations of motion, resulting in
@@
$g_{\mu\nu}=\eta_{\mu\nu}-\frac{\kappa}{2}\int\frac{\,\mathrm{d}^{D}q\delta(kq)e^{-iqx}}{(2\pi)^{D-1}}O^{-1}_{\alpha\beta\mu\nu}\mathcal{M}_{\mathrm{vertex}}^{\alpha\beta}.$ (18)
For GR, it takes the form $O^{-1}_{\alpha\beta\mu\nu}=\frac{I_{\alpha\beta\mu\nu}-\frac{1}{D-2}\eta_{\mu\nu}\eta_{\alpha\beta}}{q^{2}+i\epsilon}$, with the corresponding Schwarzschild-Tangherlini solution already derived in Ref.Khalaf and Telem (2024). Nevertheless, we focus on super-renormalizable gravity, where the action is given by Eq.(1). For simplicity, we examine a particular class of form factors, which can be chosen as
$\left\{\begin{split}&H_{0}(\Box)=H_{2}(\Box)=-\frac{\Box}{M^{2}_{\ast}}\Rightarrow F_{1}(\Box)=-\frac{F_{2}(\Box)}{2}=\frac{e^{-\frac{\Box}{M^{2}_{\ast}}}-1}{\Box},\\ &\omega(\Box)=e^{-\frac{\Box}{M^{2}_{\ast}}}.\end{split}\right$. (19)
It is straightforward to show that this choice of form factor simplifies the propagator Eq.(8) to the following form
$\begin{split}O^{-1}_{\mu\nu\alpha\beta}&=\frac{e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}I_{\mu\nu\alpha\beta}-\frac{4e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{(D-2)q^{2}}T_{\mu\nu\alpha\beta}+\frac{2e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{(D-2)q^{2}}C_{\mu\nu\alpha\beta}\\ &+\frac{(2\xi-2)e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}J_{\mu\nu\alpha\beta}+\frac{e^{-\frac{q^{2}}{M^{2}_{\ast}}}}{q^{2}}\left(\frac{D-2}{D-1}-\frac{1}{(D-2)(D-1)}-\frac{3\xi}{2}\right)K_{\mu\nu\alpha\beta}.\end{split}$ (20)
@@
Subsequently, based on the Feynman rules in Eq.(15), it can be demonstrated that the tree-level scattering amplitude in classical limit is given by $\mathcal{M}^{\mu\nu}_{\text{tree}}=-\kappa m^{2}\eta^{||}_{\mu\nu}$. Based on the coupling between the propagator and the scattering amplitudes (18), we calculate the first-order correction to the metric tensor as
$\left\{\begin{split}&h^{(G_{N})}_{\alpha\beta}=T_{1}\eta^{||}_{\alpha\beta}+T_{2}\eta^{\bot}_{\alpha\beta},\\ &T_{1}\equiv\frac{m\kappa^{2}}{2}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}\left(\frac{(D-3)e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}}{(D-2)q^{2}_{\bot}}\right)e^{-iq_{\bot}x_{\bot}},\\ &T_{2}\equiv-\frac{m\kappa^{2}}{2}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}\left(\frac{e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}}{(D-2)q^{2}_{\bot}}\right)e^{-iq_{\bot}x_{\bot}},\end{split}\right$. (21)
where we have used the $\delta$ function to express $q^{\mu}$ and $x^{\mu}$ in terms of $q^{\mu}_{\bot}$ and $x^{\mu}_{\bot}$ (Since $q^{\mu}=q^{\mu}_{||}+q^{\mu}_{\bot},x^{\mu}=x^{\mu}_{||}+x^{\mu}_{\bot}$). Observe that the integrand includes an exponential function, and we apply the following formula
$\begin{split}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}\left(\frac{e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}}{q^{2}_{\bot}}\right)e^{-iq_{\bot}x_{\bot}}=-\frac{1}{4\pi^{\frac{D-1}{2}}r^{D-3}}\left[\Gamma\left(\frac{D-3}{2}\right)-\Gamma\left(\frac{D-3}{2},\frac{1}{4}M^{2}_{\ast}r^{2}\right)\right],\end{split}$ (22)
where $\Gamma\left(\frac{D-3}{2},\frac{1}{4}M^{2}_{\ast}r^{2}\right)$ represents the upper incomplete gamma function of order $\frac{D-3}{2}$, and the variable $r$, defined as $r^{2}=-x^{2}_{\bot}$ in the reference frame of $k_{\mu}$. By substituting Eq.(22) into Eq.(LABEL:21), we ultimately obtain
@@
In particular, in the case of $D=4$, the Newtonian potential can be derived as
$\begin{split}\Phi^{(G_{N})}(r)=-\frac{G_{N}m}{r}\text{erf}\left(\frac{M_{\ast}r}{2}\right),\end{split}$ (25)
where the error function, denoted as $\text{erf}(x)$, is defined by $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}\,dt$. This result aligns with earlier work in Biswas et al. (2012); Burzillà et al. (2021) obtained with linear perturbation. The error function encodes the nonlocal smearing of the gravitational interaction: at large distances $\lim_{x\to\infty}\text{erf}(x)=1$, the potential reduces to the standard Newtonian form $\frac{G_{N}m}{r}$, recovering GR in the infrared. In the short-range regime, however, the potential remains perfectly regular: $\Phi^{(G_{N})}(r\to 0)=-\frac{G_{N}mM_{*}}{\sqrt{\pi}}$. The weak nonlocality regulates the ultraviolet behavior of gravity, defusing the classical singularity at tree-level.
IV The Metric correction from One-Loop Scattering Amplitudes
In this section, we will calculate the second-order correction to the metric in the classical limit, stemming from the one-loop scattering amplitude. The $(G_{N})^{2}$ contribution to the metric is generated by the triangle one-loop diagram depicted in Fig.2. Other one-loop topologies, devoid of such non-analyticity, do not contribute to the classical potential, thereby establishing the triangle diagram as the sole contributor to the $\mathcal{O}(G_{N}^{2})$ singularity resolution Bjerrum-Bohr et al. (2018).
@@ IV The Metric correction from One-Loop Scattering Amplitudes
where $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}$ and $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}$ represent the interaction terms involving the three-graviton vertices from GR and nonlocal terms, respectively. In addition, we introduce the tensor $f_{\alpha\beta}\equiv\frac{D-3}{D-2}\eta^{||}_{\alpha\beta}-\frac{1}{D-2}\eta^{\bot}_{\alpha\beta}$, which represents the scalar-graviton vertices contracted with the graviton propagator. These tensors encapsulate the tensor structure of the $\phi^{2}h$ vertices after they are contracted with the graviton propagator. We employ two important simplifications, which arise within the classical limit. Firstly, we neglect the factors of $l$ and $q$ in the scalar-graviton vertex. Secondly, we omit the momentum-dependent part of the graviton propagator. The first simplification is analogous to the treatment of tree-level diagram, where graviton momenta are considered negligible relative to the scalar momenta. The second simplification, involving the momentum dependence of the propagator, is less trivial, as graviton momenta appear in both the numerator and denominator of the propagator expression. This can be confirmed by examining the corresponding integrals, though it is also clear from an intuitive perspective. Since scalar-graviton vertices represent energy-momentum and are conserved, they vanish when contracted with the momentum from the propagator.
Furthermore, focusing on the tensor part of the integral involving the three-graviton vertex from the GR contribution $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}\right)$ and two factors of $f_{\alpha\beta}$, we express the amplitude from the GR part as
$\begin{split}i\mathcal{M}^{\mu\nu}_{\text{GR}}=-2m^{4}\kappa^{3}f_{\alpha\beta}f_{\gamma\delta}\left(U^{\mu\nu\alpha\beta\rho\gamma\delta\sigma}_{\text{GR}}\left(I^{\text{GR}}_{\rho\sigma}+I^{\text{GR}}_{\rho}q_{\bot\sigma}\right)+U^{\alpha\beta\gamma\delta\rho\mu\nu\sigma}_{\text{GR}}q_{\bot\rho}q_{\bot\sigma}I^{\text{GR}}\right),\end{split}$ (27)
where $U$-tensor is defined in Appendix A, and the quantities $I^{\text{GR}}_{\rho\sigma}$, $I^{\text{GR}}_{\rho}$, and $I^{\text{GR}}$ are given by
$\begin{split}I^{\text{GR}}_{a_{1}...a_{n}}\equiv\int\frac{{\,\mathrm{d}^{D}l}}{(2\pi)^{D}}\frac{e^{-\frac{l^{2}}{M^{2}_{\ast}}}e^{-\frac{\left(l+q_{\bot}\right)^{2}}{M^{2}_{\ast}}}l_{a_{1}}...l_{a_{n}}}{l^{2}(l+q_{\bot})^{2}\left((l+k)^{2}-m^{2}+i\epsilon\right)}.\end{split}$ (28)
Using the integral formulas provided in Appendix B, we can show that the term $I^{\text{GR}}_{\rho\sigma}+I^{\text{GR}}_{\rho}q_{\bot\sigma}$ in Eq.(27) can be expressed as
$\begin{split}I^{\text{GR}}_{\rho\sigma}+I^{\text{GR}}_{\rho}q_{\bot\sigma}=\frac{iq^{2}_{\bot}N_{D-1}}{16\left(D-2\right)m}\left(\left(D-3\right)\frac{q_{\bot\rho}q_{\bot\sigma}}{q^{2}_{\bot}}+\eta^{\bot}_{\rho\sigma}\right)+\frac{iN_{D}}{4m^{2}}\left(k_{\rho}q_{\bot\sigma}-q_{\bot\rho}k_{\sigma}\right),\end{split}$ (29)
here the integrals for $N_{D-1}$ and $N_{D}$ are also defined in Appendix B. By applying the results above and the vertex rule, we obtain the final contribution from the GR part as
$\begin{split}i\mathcal{M}^{\mu\nu}_{\text{GR}}=\frac{im^{3}\kappa^{3}q^{2}_{\bot}N_{D-1}}{2}\left(\frac{7D-25}{8(D-2)}\eta_{||}^{\mu\nu}-\frac{5D^{2}-18D+17}{8(D-2)}\left(\eta^{\mu\nu}_{\bot}-\frac{q_{\bot\rho}q_{\bot\sigma}}{q^{2}_{\bot}}\right)\right).\end{split}$ (30)
Subsequently, using the vertex rule for nonlocal terms provided in Appendix A (denoted as $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}$), we obtain
@@
where we define three types of integrals, which primarily differ due to the contributions from the form factor $F_{1}(\Box)$. These integrals can be written as
[…]
Similarly, based on the definition of the full amplitude contribution from nonlocal terms, we obtain
[…]
the integrals $N^{(i)}_{D},N^{(i)}_{D-1},i=1,2,3$ are given by Appendix B. From the above expression, it is evident that the contribution from the nonlocal terms introduces additional term involving $k^{\mu}q^{\nu}_{\bot}+k^{\nu}q^{\mu}_{\bot}$ compared to the GR. This indicates that such term contribute to correction in the spacetime cross terms of the metric. Furthermore, we find that the nonlocal terms also include an extra $q^{2}_{\bot}$ term in the numerator compared to the GR term. This suggests that, in the classical limit, the GR contribution will dominate. It also implies that the theory has a certain range of feasible scales within which it is valid.
@@
Ultimately, by substituting Eq.(30) and Eq.(LABEL:34) into Eq.(18), we obtain the corrections to the metric and Newtonian potential up to second order in $(G_{N})^{2}$ within the one-loop approximation. These corrections can be expressed as
@@
$\left\{\begin{split}&h^{(G_{N})^{2}}_{\mu\nu}\equiv h^{(G_{N})^{2}\text{GR}}_{\mu\nu}+h^{(G_{N})^{2}\text{Nonlocal}}_{\mu\nu},\\ &h^{(G_{N})^{2}\text{GR}}_{\mu\nu}=-\frac{\kappa^{4}m^{2}}{2}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}e^{-iq_{\bot}x_{\bot}}e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}N_{D-1}\left[\frac{183D+5D^{3}-61D^{2}-151}{16(D-2)^{2}}\eta^{||}_{\mu\nu}\right.\\ &\left.+\frac{25-7D}{16(D-2)^{2}}\eta^{\bot}_{\mu\nu}+\frac{(17+D)(5D-18)}{16(D-2)}\frac{q_{\mu}q_{\nu}}{q^{2}_{\bot}}\right],\\ &h^{(G_{N})^{2}\text{Nonlocal}}_{\mu\nu}=-\frac{\kappa^{4}m^{2}}{2}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}e^{-iq_{\bot}x_{\bot}}e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}\left[\frac{3N^{(1)}_{D}\left(16-16D+3D^{2}\right)}{16m(D-1)(D-2)^{2}}\left(k_{\mu}q_{\bot\nu}+k_{\nu}q_{\bot\mu}\right)\right.\\ &\left.-\frac{N^{(2)}_{D}(D-1)}{2m(D-2)^{2}}\left(k_{\mu}q_{\bot\nu}+k_{\nu}q_{\bot\mu}\right)-\frac{N^{(3)}_{D}\left(8D^{3}-43D^{2}+72D-32\right)}{16m(D-1)(D-2)^{2}}\left(k_{\mu}q_{\bot\nu}+k_{\nu}q_{\bot\mu}\right)\right.\\ &\left.-q^{2}_{\bot}N^{{(1)}}_{D-1}\left(\frac{D^{2}-5D+8}{4(D-2)^{3}}\eta^{||}_{\mu\nu}-\frac{4D-19}{16(D-2)^{3}}\eta^{\bot}_{\mu\nu}-\frac{3(D-1)}{16(D-2)^{3}}\frac{q_{\bot\mu}q_{\bot\nu}}{q^{2}_{\bot}}\right)\right.\\ &\left.-q^{2}_{\bot}N^{{(3)}}_{D-1}\left(\frac{D^{2}-5D+8}{4(D-2)^{3}}\eta^{||}_{\mu\nu}-\frac{4D-19}{16(D-2)^{3}}\eta^{\bot}_{\mu\nu}-\frac{13-5D}{16(D-2)^{3}}\frac{q_{\bot\mu}q_{\bot\nu}}{q^{2}_{\bot}}\right)\right.\\ &\left.-q^{2}_{\bot}N^{{(2)}}_{D-1}\left(\frac{76+13D+10D^{2}}{8(D-2)^{3}}\eta^{||}_{\mu\nu}-\frac{7D-14D^{2}-101}{8(D-2)^{3}}\eta^{\bot}_{\mu\nu}-\frac{107-121D+46D^{2}-6D^{3}}{8(D-2)^{3}}\frac{q_{\bot\mu}q_{\bot\nu}}{q^{2}_{\bot}}\right)\right],\end{split}\right$. (34)
with $(G_{N})^{2}$ order Newtonian potential
$\begin{split}\Phi^{(G_{N})^{2}}(r)&=-\frac{\kappa^{4}m^{2}}{4}\int\frac{\,\mathrm{d}^{D-1}q_{\bot}}{(2\pi)^{D-1}}e^{-iq_{\bot}x_{\bot}}e^{-\frac{q^{2}_{\bot}}{M^{2}_{\ast}}}\left[N_{D-1}\frac{183D+5D^{3}-61D^{2}-151}{16(D-2)^{2}}\right.\\ &\left.-q^{2}_{\bot}\left(N^{(1)}_{D-1}+N^{(3)}_{D-1}\right)\left(\frac{D^{2}-5D+8}{4(D-2)^{3}}\right)-q^{2}_{\bot}N^{(2)}_{D-1}\left(\frac{76+13D+10D^{2}}{8(D-2)^{3}}\right)\right].\end{split}$ (35)
@@
Based on the above equation, we find that the Newtonian potential at order $\mathcal{O}(G_{N}^{2})$, given in Eq. (35), is structurally governed by the spacetime dimension $D$ and four loop integrals $N_{D-1},N^{(i)}_{D-1}$ (with $i=1,2,3$), whose explicit forms are provided in Appendix B. These integrals do not admit a simple analytic expression, motivating a numerical treatment in the physically relevant case $D=4$. As displayed in Fig. 3, the computed potential is plotted against the dimensionless radial coordinate, the one-loop correction grows smoothly with distance, matching expected long-range behavior. Most notably, the potential remains completely regular as $r\to 0$, offering direct numerical evidence that the mechanism of singularity resolution, enforced by weak nonlocality, persists at the quantum level.